# The Complexity of Resilience Problems via Valued Constraint Satisfaction

### v5

@@ 5. Hardness from pp-Constructions
Let $C_{1}$, $C_{2}$, $C_{3}$ be the domains of $\Gamma_{1}$, $\Gamma_{2}$, and $\Gamma_{3}$, respectively. If $a\in C_{1}^{k}$ and $c\in C_{3}^{k}$, for some $k\in{\mathbb{N}}$, then define
$\omega_{3}({\mathscr{S}}_{a,c}):=\sum_{b\in C_{2}^{k}}\omega_{1}({\mathscr{S}}_{a,b})\omega_{2}({\mathscr{S}}_{b,c}).$
Note that on sets of this form, i.e., on basic open sets in $C_{3}^{C_{1}}$, $\omega_{3}$ is countably additive. Since our basis of open sets is closed under intersection, this definition extends uniquely to all of $B(C_{3}^{C_{1}})$ by Dynkin’s $\pi$-$\lambda$ theorem. ∎
@@
The following was shown for valued structures over finite domains in (Butti, 2022, Proposition 8.4).
@@
We say that two valued $\tau$-structures $\Gamma$ and $\Delta$ are fractionally homomorphically equivalent if there exists a fractional homomorphisms from $\Gamma$ to $\Delta$ and from $\Delta$ to $\Gamma$. Clearly, fractional homomorphic equivalence is indeed an equivalence relation on valued structures of the same signature.
@@
Let $\Gamma$ and $\Delta$ be valued $\tau$-structures with oligomorphic automorphism groups that are fractionally homomorphically equivalent. Let $\omega$ be a fractional homomorphism from $\Delta$ to $\Gamma$. Let $R\in\langle\Delta\rangle$ be of arity $k$ and $R^{\prime}\in\langle\Gamma\rangle$ be the valued relation obtained when the expression for $R$ in $\Delta$ is interpreted over $\Gamma$. Then for every $a\in D^{k}$,
@@
Recall that by Lemma 5.5, the expected values from (5) exist. By the definition of a fractional homomorphism, (5) holds for every pair $(R,R^{\prime})=(S^{\Delta},S^{\Gamma})$ where $S\in\tau$. Clearly, the same is true for $R=R_{\emptyset}$. To see that (5) holds for $R=R_{=}$, let $a\in D^{2}$. Note that either $a_{1}=a_{2}$ in which case $f(a_{1})=f(a_{2})$ for every $f\in C^{D}$, and hence both sides of (5) are equal to 00, or $a_{1}\neq a_{2}$, in which case $R(a)=\infty$ and (5) is again satisfied.
We will show that every valued relation $R$ obtained from a valued relation satisfying (5) by a single operator from Definition 4.1 satisfies (5); the general statement then follows by induction. This is clear for valued relations $R$ obtained by non-negative scaling and addition of constants, since these operations preserve (5) by the linearity of expectation. The assumption that $\Gamma$ and $\Delta$ are fractionally homomorphically equivalent (rather than the existence of $\omega$) is needed only for the operator $\operatorname{Opt}$.
Let $\phi(x_{1},\dots,x_{k},y_{1},\dots,y_{n})$ be a $\tau$-expression. Let $R$ be the $k$-ary valued relation defined by $R(x)=\inf_{y\in D^{n}}\phi^{\Delta}(x,y)$ for every $x\in D^{k}$ Since $\phi$ is a $\tau$-expression, there are $R_{i}\in\tau$ such that
@@
Note that the hardness proofs in Examples 4.7 and 4.8 are special cases of Corollary 5.13. A more involved example uses the relational structure with a hard CSP introduced below and is shown in Example 8.25.
@@
Let $\Gamma$ be a valued structure with a finite signature and oligomorphic automorphism group such that ($(\{0,1\};\operatorname{OIT})$ has a pp-construction in $\Gamma$. Then $\operatorname{VCSP}(\Gamma)$ is NP-hard.
Since Corollary 5.14 is our typical tool to prove hardness, we are interested in pp-constructing relational structures in valued structures. If $\Gamma$ is a valued structure on a domain $C$, then we call the relational structure on the domain $C$ whose relations are all classical relations from $\langle\Gamma\rangle$ the underlying crisp structure of $\Gamma$. The following proposition first appeared in (Bodirsky et al., 2024a) and shows that for hardness proofs by Corollary 5.14, it is equivalent to check pp-constructability in the underlying crisp structure.
@@
Let $\Gamma$ be a valued structure and let ${\mathfrak{B}}$ be a relational structure, both with countable domains. Then $\Gamma$ pp-constructs ${\mathfrak{B}}$ if and only if the underlying crisp structure of $\Gamma$ pp-constructs ${\mathfrak{B}}$.
@@
The relation of pp-constructibility on the class of countable valued structures is transitive.
@@
Clearly, a pp-power of a pp-power is again a pp-power, and fractional homomorphic equivalence is transitive by Lemma 5.6. We are therefore left to prove that if $\Gamma$ and $\Delta$ are valued structures such that $\Delta$ is a $d$-dimensional pp-power of $\Gamma$, and if $\Gamma^{\prime}$ is fractionally homomorphically equivalent to $\Gamma$ via fractional homomorphisms $\omega_{1}\colon\Gamma\to\Gamma^{\prime}$ and $\omega_{2}\colon\Gamma^{\prime}\to\Gamma$, then $\Delta$ also has a pp-construction in $\Gamma^{\prime}$.
Let $C$ and $C^{\prime}$ be the domains of $\Gamma$ and $\Gamma^{\prime}$, respectively. Take the $\tau$-expressions that define the valued relations of $\Delta$ over $\Gamma$, and interpret them over $\Gamma^{\prime}$ instead of $\Gamma$; let $\Delta^{\prime}$ be the resulting valued structure. Note that $\Delta^{\prime}$ is a $d$-dimensional pp-power of $\Gamma^{\prime}$. For a map $f\colon\Gamma\to\Gamma^{\prime}$, let $\tilde{f}\colon\Delta\to\Delta^{\prime}$ be given by $(x_{1},\dots,x_{d})\mapsto(f(x_{1}),\dots,f(x_{d}))$. Then for all $S\in B((C^{\prime})^{C})$ we define
@@ 6. Fractional Polymorphisms
In this section we introduce fractional polymorphisms of valued structures; they are an important tool for formulating tractability results and complexity classifications of VCSPs. For valued structures with a finite domain, our definition specialises to the established notion of a fractional polymorphism which has been used to study the complexity of VCSPs for valued structures over finite domains (see, e.g. (Thapper and Živný, 2013)). Our approach is different from the one of Schneider and Viola (Viola, 2020; Schneider and Viola, 2022) and Viola and Živný (Viola and Živný, 2021) in that we work with arbitrary probability spaces instead of distributions with finite support or countable additivity property. As we will see in Section 7, fractional polymorphisms can be used to give sufficient conditions for tractability of $\operatorname{VCSP}$s of certain valued structures with oligomorphic automorphism groups. This justifies our more general notion of a fractional polymorphism, as it might provide a tractability proof for more problems.
@@
If the reference to $C$ is clear, we occasionally omit the subscript $C$. We often use $\omega$ for both the entire fractional operation and for the map $\omega\colon B({\mathscr{O}}_{C}^{(\ell)})\to[0,1]$.
@@
Note that (10) has the interpretation that the expected value of $R(f(a^{1},\dots,a^{\ell}))$ is at most the average of the values $R(a^{1}),\dots$, $R(a^{\ell})$. Also note that if $R$ is a classical relation improved by a fractional operation $\omega$ and $\omega(f)>0$ for $f\in{\mathscr{O}}^{(\ell)}$, then $f$ must preserve $R$ in the usual sense. It follows from Lemma 5.5 that if $\operatorname{Aut}(C;R)$ is oligomorphic, then $E_{\omega}[f\mapsto R(f(a^{1},\dots,a^{\ell}))]$ always exists and is greater than $\infty$- ∞.
@@
Let $<$ be the binary relation on $\{0,1\}$ and $\Gamma_{<}$ the valued structure from Example 2.2. By definition, $\operatorname{Opt}(<)\in\langle\Gamma_{<}\rangle$. Denote the minimum operation on $\{0,1\}$ by $\min$ and let $\omega$ be a binary fractional operation defined by $\omega(\min)=1$. Note that $\omega\in\operatorname{fPol}(\{0,1\};\operatorname{Opt}(<))$. However,
$<\left(\min\left(\begin{pmatrix}0\\ 1\end{pmatrix},\begin{pmatrix}0\\ 0\end{pmatrix}\right)\right)=\,{<}(0,0)=1,$
[…]

### v6

@@ 5. Hardness from pp-Constructions
Since we are working with Borel $\sigma$-algebras, it is enough to prove that $\circ$ is continuous. Therefore, it is enough to verify that preimages of basic open sets in $C_{3}^{C_{1}}$ are open. Let $a\in C_{1}^{k}$ and $c\in C_{3}^{k}$ for some $k\in\mathbb{N}$. Then
$\circ^{-1}(\mathscr{S}_{a,c})=\bigcup_{b\in C_{2}^{k}}(\mathscr{S}_{b,c}\times\mathscr{S}_{a,b}),$
which is an open set in $C_{3}^{C_{2}}\times C_{2}^{C_{1}}$. It follows that $\circ$ is continous and therefore measurable. ∎
Definition 5.8 (composition of fractional maps).
Let $C_{1},C_{2},C_{3}$ be countable sets. Let $\omega_{1}$ be a fractional map from $C_{1}$ to $C_{2}$ and $\omega_{2}$ be a fractional map from $C_{2}$ to $C_{3}$. Then $\omega_{2}\circ\omega_{1}\colon\mathcal{B}(C_{3}^{C_{1}})\to[0,1]$ is defined by
$\omega_{2}\circ\omega_{1}(S):=\omega_{2}\otimes\omega_{1}(\{(g,f)\mid f\in C_{2}^{C_{1}},g\in C_{3}^{C_{2}},g\circ f\in S\}).$
Note that the set $\{(g,f)\mid f\in C_{2}^{C_{1}},g\in C_{3}^{C_{2}},g\circ f\in S\}$ is the preimage of $S\in\mathcal{B}(C_{3}^{C_{1}})$ under $\circ$. The map $\circ\colon C_{3}^{C_{2}}\times C_{2}^{C_{1}}\to C_{3}^{C_{1}}$ is measurable by Lemma 5.7, and therefore the preimage of $S$ under $\circ$ lies in ${\mathcal{B}}(C_{3}^{C_{2}}\times C_{2}^{C_{1}})$. Since $C_{3}^{C_{2}}$ and $C_{2}^{C_{1}}$ are second-countable topological spaces, ${\mathcal{B}}(C_{3}^{C_{2}})\otimes{\mathcal{B}}(C_{2}^{C_{1}})={\mathcal{B}}(C_{3}^{C_{2}}\times C_{2}^{C_{1}})$ by Lemma 5.6. Therefore, the set $\{(g,f)\mid f\in C_{2}^{C_{1}},g\in C_{3}^{C_{2}},g\circ f\in S\}$ lies in ${\mathcal{B}}(C_{3}^{C_{2}})\otimes{\mathcal{B}}(C_{2}^{C_{1}})$, and hence $\omega_{2}\circ\omega_{1}$ is well-defined. We proceed by proving several natural properties for the operation $\circ$.
Proposition 5.9.
Let $C_{1},C_{2},C_{3}$ be countable sets. Let $\omega_{1}$ be a fractional map from $C_{1}$ to $C_{2}$ and let $\omega_{2}$ be a fractional map from $C_{2}$ to $C_{3}$. Then the following statements hold.
  1. (1)
    $\omega_{2}\circ\omega_{1}$ is a fractional map from $C_{1}$ to $C_{3}$.
  2. (2)
    For every $k\in\mathbb{N}$, $a\in C_{1}^{k}$, and $c\in C_{3}^{k}$,
    $\omega_{2}\circ\omega_{1}(\mathscr{S}_{a,c})=\sum_{b\in C_{2}^{k}}\omega_{1}({\mathscr{S}}_{a,b})\omega_{2}({\mathscr{S}}_{b,c}).$
  3. (3)
    The above property uniquely determines the fractional map $\omega_{2}\circ\omega_{1}$.
  4. (4)
    The operation $\circ$ of composition of fractional maps is associative.
  5. (5)
    If $\Gamma_{1},\Gamma_{2}$, and $\Gamma_{3}$ are valued $\tau$-structures, $\omega_{1}$ is a fractional homomorphism from $\Gamma_{1}$ to $\Gamma_{2}$, and $\omega_{2}$ is a fractional homomorphism from $\Gamma_{2}$ to $\Gamma_{3}$, then $\omega_{2}\circ\omega_{1}$ is a fractional homomorphism from $\Gamma_{1}$ to $\Gamma_{3}$.
Proof.
To prove item (1), note that $\omega_{2}\circ\omega_{1}$ is the pushforward of $\omega_{2}\otimes\omega_{1}$ by the measurable map $\circ\colon C_{3}^{C_{2}}\times C_{2}^{C_{1}}\to C_{3}^{C_{1}}$; see, e.g., (Tao, 2011, Exercise 1.4.38) for the definition. This construction is known to yield a probability distribution (see again (Tao, 2011, Exercise 1.4.38)), in other words, a fractional map.
To show item (2), let $k\in\mathbb{N}$, $a\in C_{1}^{k}$, and $c\in C_{3}^{k}$. Then
$\omega_{2}\circ\omega_{1}(\mathscr{S}_{a,c})$ $=\omega_{2}\otimes\omega_{1}(\{(g,f)\mid f\in C_{2}^{C_{1}},g\in C_{3}^{C_{2}},g\circ f\in\mathscr{S}_{a,c}\})$
$=\omega_{2}\otimes\omega_{1}(\{(g,f)\mid f\in C_{2}^{C_{1}},g\in C_{3}^{C_{2}},\exists b\in C_{2}^{k}:f(a)=b\wedge g(b)=c\})$
$=\omega_{2}\otimes\omega_{1}\left(\bigcup_{b\in C_{2}^{k}}\{(g,f)\mid f\in C_{2}^{C_{1}},g\in C_{3}^{C_{2}},f(a)=b\wedge g(b)=c\}\right)$
$=\sum_{b\in C_{2}^{k}}\omega_{2}\otimes\omega_{1}\left(\mathscr{S}_{b,c}\times\mathscr{S}_{a,b}\right)$
$=\sum_{b\in C_{2}^{k}}\omega_{2}(\mathscr{S}_{b,c})\cdot\omega_{1}(\mathscr{S}_{a,b}).$
Item (3) is a consequence of Dynkin’s $\pi$-$\lambda$ theorem (see, e.g., (Aliprantis and Border, 2006, Lemma 4.11)), because basic open sets form a $\pi$-system (a non-empty system closed under intersections).
To show item (4), let $\omega_{3}$ be a fractional map from $C_{3}$ to a countable set $C_{4}$. Let $S\in\mathcal{B}(C_{4}^{C_{1}})$. We need to show $\omega_{3}\circ(\omega_{2}\circ\omega_{1})(S)=(\omega_{3}\circ\omega_{2})\circ\omega_{1}(S)$. Note that it is enough to verify this property for basic open sets $S$, since then, again by Dynkin’s $\pi$-$\lambda$ theorem, $\omega_{3}\circ(\omega_{2}\circ\omega_{1})=(\omega_{3}\circ\omega_{2})\circ\omega_{1}$. Let $k\in\mathbb{N}$, $a\in C_{1}^{k}$ and $d\in C_{4}^{k}$. Then
$\omega_{3}\circ(\omega_{2}\circ\omega_{1})(\mathscr{S}_{a,d})$ $=\omega_{3}\otimes(\omega_{2}\circ\omega_{1})\left(\left\{(h,h^{\prime})\mid h\in C_{4}^{C_{3}},h^{\prime}\in C_{3}^{C_{1}},h\circ h^{\prime}(a)=d\right\}\right)$
$=\omega_{3}\otimes(\omega_{2}\circ\omega_{1})\left(\bigcup_{c\in C_{3}^{k}}\mathscr{S}_{c,d}\times\mathscr{S}_{a,c}\right)$
$=\sum_{c\in C_{3}^{k}}\omega_{3}(\mathscr{S}_{c,d})\cdot(\omega_{2}\circ\omega_{1})(\mathscr{S}_{a,c})$
$=\sum_{c\in C_{3}^{k}}\omega_{3}(\mathscr{S}_{c,d})\cdot(\omega_{2}\otimes\omega_{1})\left(\left\{(g,f)\mid g\in C_{3}^{C_{2}},f\in C_{2}^{C_{1}},g\circ f(a)=c\right\}\right)$
$=\sum_{c\in C_{3}^{k}}\omega_{3}(\mathscr{S}_{c,d})\cdot(\omega_{2}\otimes\omega_{1})\left(\bigcup_{b\in C_{2}^{k}}\mathscr{S}_{b,c}\times\mathscr{S}_{a,b}\right)$
$=\sum_{c\in C_{3}^{k}}\sum_{b\in C_{2}^{k}}\omega_{3}(\mathscr{S}_{c,d})\cdot\omega_{2}(\mathscr{S}_{b,c})\cdot\omega_{1}(\mathscr{S}_{a,b})$
$=\sum_{b\in C_{2}^{k}}\omega_{1}(\mathscr{S}_{a,b})(\omega_{3}\otimes\omega_{2})\left(\bigcup_{c\in C_{3}^{k}}\mathscr{S}_{c,d}\times\mathscr{S}_{b,c}\right)$
$=\dots=(\omega_{3}\circ\omega_{2})\circ\omega_{1}(\mathscr{S}_{a,d})$
Note that the order of the summations can be exchanged by the discrete version of the Fubini-Tonelli theorem (Klenke, 2020, Theorem 14.19), because the summands are non-negative.
To prove item (5), let $R\in\tau$ be of arity $k$ and $a\in C_{1}^{k}$. Then by item (3) we have the following:
$E_{\omega_{2}\circ\omega_{1}}[h\mapsto R(h(a))]$ $=\sum_{c\in C_{3}^{k}}\omega_{2}\circ\omega_{1}(\mathscr{S}_{a,c})\cdot R(c)$
$=\sum_{c\in C_{3}^{k}}R(c)\sum_{b\in C_{2}^{k}}\omega_{2}(\mathscr{S}_{b,c})\cdot\omega_{1}(\mathscr{S}_{a,b})$
$=\sum_{b\in C_{2}^{k}}\omega_{1}(\mathscr{S}_{a,b})\sum_{c\in C_{3}^{k}}R(c)\omega_{2}(\mathscr{S}_{b,c})$
$=\sum_{b\in C_{2}^{k}}\omega_{1}(\mathscr{S}_{a,b})E_{\omega_{2}}[g\mapsto R(g(b))]$
$\leq\sum_{b\in C_{2}^{k}}\omega_{1}(\mathscr{S}_{a,b})\cdot R(b)$
$=E_{\omega_{1}}[f\mapsto R(f(a))]$
$\leq R(a),$
where the two inequalities follow from $\omega_{1}$ and $\omega_{2}$ being fractional homomorphisms. The order of the summations can again be exchanged by the Fubini-Tonelli theorem. It follows that $\omega_{2}\circ\omega_{1}$ is a fractional homomorphism. ∎
@@
The following property of fractional homomorphisms was shown for valued structures over finite domains in (Butti, 2022, Proposition 8.4).
@@
We say that two relational $\tau$-structures $\mathfrak{A}$} and $\mathfrak{B}$} are homomorphically equivalent if there is a homomorphism from $\mathfrak{A}$} to $\mathfrak{B}$} and from $\mathfrak{B}$} to $\mathfrak{A}$}. We say that two valued $\tau$-structures $\Gamma$ and $\Delta$ are fractionally homomorphically equivalent if there exists a fractional homomorphism from $\Gamma$ to $\Delta$ and from $\Delta$ to $\Gamma$. Clearly, fractional homomorphic equivalence is indeed an equivalence relation on valued structures of the same signature.
@@
Let $\Gamma$ and $\Delta$ be valued $\tau$-structures with oligomorphic automorphism groups that are fractionally homomorphically equivalent. Let $\omega$ be a fractional homomorphism from $\Delta$ to $\Gamma$. Let $R\in\langle\Delta\rangle$ be of arity $k$ and $R^{\prime}\in\langle\Gamma\rangle$ be the valued relation obtained when the pp-expression for $R$ in $\Delta$ is interpreted over $\Gamma$. Then for every $a\in D^{k}$,
@@
Recall that by Lemma 5.5, the expected values from (5) exist. By the definition of a fractional homomorphism, (5) holds for every pair $(R,R^{\prime})=(S^{\Delta},S^{\Gamma})$ where $S\in\tau$. Clearly, the same is true for $R=R_{\emptyset}$. To see that (5) holds for $R=R_{=}$, let $a\in D^{2}$. Note that either $a_{1}=a_{2}$ in which case $f(a_{1})=f(a_{2})$ for every $f\in C^{D}$, and hence both sides of (5) are equal to 0, or $a_{1}\neq a_{2}$, in which case $R(a)=\infty$ and (5) is again satisfied.
We will show that every valued relation $R$ obtained from a valued relation satisfying (5) by an application of an operator from Definition 4.2 satisfies (5); the general statement then follows by induction. This is clear for valued relations $R$ obtained by non-negative scaling and addition of constants, since these operations preserve (5) by the linearity of expectation. The assumption that $\Gamma$ and $\Delta$ are fractionally homomorphically equivalent (rather than the existence of $\omega$) is needed only for the operator $\operatorname{Opt}$.
Let $\phi(x_{1},\dots,x_{k},y_{1},\dots,y_{n})$ be a $\tau$-expression. Let $R$ be the $k$-ary valued relation defined by $R(x)=\inf_{y\in D^{n}}\phi^{\Delta}(x,y)$ for every $x\in D^{k}$ . Since $\phi$ is a $\tau$-expression, there are $R_{i}\in\tau$ such that
@@
We now have all ingredients needed to introduce the notion of a pp-construction.
@@
Note that the hardness proofs in Examples 4.6 and 4.7 are special cases of Corollary 5.16. A more involved example uses the relational structure with a hard CSP introduced below and is shown in Example 8.27.
@@
Let $\Gamma$ be a valued structure with a finite signature and oligomorphic automorphism group such that $(\{0,1\};\operatorname{OIT})$ has a pp-construction in $\Gamma$. Then $\operatorname{VCSP}(\Gamma)$ is NP-hard.
Since Corollary 5.17 is our typical tool to prove hardness, we are interested in pp-constructing relational structures in valued structures. If $\Gamma$ is a valued structure on a domain $C$, then we call the relational structure on the domain $C$ whose relations are all classical relations (i.e., with values in $\{0,\infty\}$) in $\langle\Gamma\rangle$ the underlying crisp structure of $\Gamma$. The following proposition first appeared in (Bodirsky et al., 2024a) and shows that for hardness proofs by Corollary 5.17, it is equivalent to check pp-constructability in the underlying crisp structure.
@@
Let $\Gamma$ be a valued structure and let $\mathfrak{B}$} be a relational structure, both with countable domains. Then $\Gamma$ $p$-constructs $\mathfrak{B}$} if and only if the underlying crisp structure of $\Gamma$ $p$-constructs $\mathfrak{B}$}.
@@
The relation of pp-constructability on the class of countable valued structures is transitive.
@@
Clearly, a pp-power of a pp-power is again a pp-power, and fractional homomorphic equivalence is transitive by Proposition 5.9. We are therefore left to prove that if $\Gamma$ and $\Delta$ are valued structures such that $\Delta$ is a $d$-dimensional pp-power of $\Gamma$, and if $\Gamma^{\prime}$ is fractionally homomorphically equivalent to $\Gamma$ via fractional homomorphisms $\omega_{1}\colon\Gamma\to\Gamma^{\prime}$ and $\omega_{2}\colon\Gamma^{\prime}\to\Gamma$, then $\Delta$ also has a pp-construction in $\Gamma^{\prime}$.
Let $C$ and $C^{\prime}$ be the domains of $\Gamma$ and $\Gamma^{\prime}$, respectively. Take the pp-expressions that define the valued relations of $\Delta$ over $\Gamma$, and interpret them over $\Gamma^{\prime}$ instead of $\Gamma$; let $\Delta^{\prime}$ be the resulting valued structure. Note that $\Delta^{\prime}$ is a $d$-dimensional pp-power of $\Gamma^{\prime}$. For a map $f\colon\Gamma\to\Gamma^{\prime}$, let $\tilde{f}\colon\Delta\to\Delta^{\prime}$ be given by $(x_{1},\dots,x_{d})\mapsto(f(x_{1}),\dots,f(x_{d}))$. Then for all $S\in B((C^{\prime})^{C})$ we define
@@ 6. Fractional Polymorphisms
In this section we introduce fractional polymorphisms of valued structures; they are an important tool for formulating tractability results and complexity classifications of VCSPs. For valued structures with a finite domain, our definition specializes to the established notion of a fractional polymorphism which has been used to study the complexity of VCSPs for valued structures over finite domains (see, e.g. (Thapper and Živný, 2013)). Our approach is different from the one of Schneider and Viola (Viola, 2020; Schneider and Viola, 2022) and Viola and Živný (Viola and Živný, 2021) in that we work with arbitrary probability spaces instead of distributions with finite support or countable additivity property. As we will see in Section 7, fractional polymorphisms can be used to give sufficient conditions for tractability of $\operatorname{VCSP}$s of certain valued structures with oligomorphic automorphism groups. This justifies our more general notion of a fractional polymorphism, as it might provide a tractability proof for more problems.
@@
If the reference to $C$ is clear, we occasionally omit the subscript $C$. We often use $\omega$ for both the entire fractional operation and for the map $\omega\colon B({\mathscr{O}}_{C}^{(\ell)})\to[0,1]$. Note that fractional operations are simply fractional maps where $D=C^{\ell}$ (see Definition 5.3).
@@
Note that (10) has the interpretation that the expected value of $R(f(a^{1},\dots,a^{\ell}))$ is at most the average of the values $R(a^{1}),\dots$, $R(a^{\ell})$. Also note that if $R$ is a classical relation improved by a fractional operation $\omega$ and $\omega(f)>0$ for $f\in{\mathscr{O}}^{(\ell)}$, then $f$ must preserve $R$ in the usual sense. It follows from Lemma 5.5 that if $\operatorname{Aut}(C;R)$ is oligomorphic, then $E_{\omega}[f\mapsto R(f(a^{1},\dots,a^{\ell}))]$ always exists and is greater than -$\infty$.
@@
Let << be the binary relation on $\{0,1\}$ and $\Gamma_{<}$ the valued structure from Example 2.1. By definition, $\operatorname{Opt}(<)\in\langle\Gamma_{<}\rangle$. Denote the minimum operation on $\{0,1\}$ by $\min$ and let $\omega$ be a binary fractional operation defined by $\omega(\min)=1$. Note that $\omega\in\operatorname{fPol}(\{0,1\};\operatorname{Opt}(<))$. However,
$\left(\min\left(\begin{pmatrix}0\\ 1\end{pmatrix},\begin{pmatrix}0\\ 0\end{pmatrix}\right)\right)=\,{<}(0,0)=1,$
while $(1/2)\cdot{<}(0,1)+(1/2)\cdot{<}(0,0)=1/2$. This shows that $\omega$ does not improve << and hence $\not\in\langle(\{0,1\};\operatorname{Opt}(<))\rangle$ by Lemma 6.9.